### v1

# CoSteer: Collaborative Decoding-Time Personalization via Local Delta Steering

Abstract
Personalized text generation has become crucial for adapting language models to diverse and evolving users’ personal context across cultural, temporal, and contextual dimensions. While existing methods often rely on centralized fine-tuning or static preference alignment, they struggle to achieve real-time adaptation under resource constraints inherent to personal devices. This limitation creates a dilemma: large cloud-based models lack access to localized user-specific information, while small on-device models cannot match the generation quality of their cloud counterparts. To address this dichotomy, we present CoSteer, a novel collaborative framework that enables decoding-time personalization through localized delta steering. Our key insight lies in leveraging the logits difference between personal context-aware and -agnostic outputs from local small models as steering signals for cloud-based LLMs. Specifically, we formulate token-level optimization as an online learning problem, where local delta vectors dynamically adjust the remote LLM’s logits within the on-device environment. This approach preserves privacy by transmitting only the final steered tokens rather than raw data or intermediate vectors, while maintaining cloud-based LLMs’ general capabilities without fine-tuning. Through comprehensive experiments on various personalized generation tasks, we demonstrate that CoSteer effectively assists LLMs in generating personalized content by leveraging locally stored user profiles and histories, ensuring privacy preservation through on-device data processing while maintaining acceptable computational overhead.
11footnotetext: Equal Contribution.11footnotetext: Corresponding authors.
1 Introduction
The rapid development of large language models (LLMs) has significantly enhanced natural language processing, enabling these models to understand context and generate coherent text effectively [1, 2]. This progress has sparked interest in personalization, where AI systems move beyond generic content to tailor interactions based on individual user profiles. Personalized generation refers to creating content customized through analysis of user-specific attributes like linguistic patterns, interaction histories, and contextual preferences [3]. This approach enables user-specific outputs that maintain contextual relevance, with applications in personalized recommender systems [4, 5, 6], adaptive dialogue agents [7], and customized content creation platforms [8].
Existing personalized generation approaches primarily fall into two paradigms. The first involves training-based methods, which leverage user data to tailor models using two main strategies. Parameter-efficient architectures, such as PLoRA [9], employ a shared LoRA module across users to strike a balance between personalization and resource efficiency. In contrast, individualized adapters, as seen in UserAdapter [10], optimize for per-user customization, albeit at a higher computational cost. Additionally, multi-objective reinforcement learning frameworks [11, 12] redefine personalization as the alignment of generic model capabilities with user-specific patterns, achieved through reward shaping. The second paradigm comprises tuning-free methods, which avoid model updates by adapting to user context during inference. This approach is further divided into two techniques. Prompt engineering implicitly encodes user profiles through curated prompts. For example, Cue-CoT [13] integrates user-specific reasoning into system prompts, while PAG [14] retrieves historical interactions to construct context-aware prompts. On the other hand, inference-time optimization techniques, exemplified by PAD [15] and Drift [16], actively adjust token generation probabilities by manipulating logit differences. Unlike prompt-based methods that inject semantic context, these techniques constrain output distributions to better align with user preferences.
However, both training-based and tuning-free methods face critical dilemmas in balancing privacy and quality when deployed on resource-constrained personal devices. Training-based methods require computational resources that may not be available locally, and cloud-based training risks exposing user data. Meanwhile, tuning free methods that transmit raw personal context to cloud-based LLMs for context-aware adaptation also risk exposing sensitive information, whereas relying solely on local SLMs leads to degraded output quality. Therefore, there is a pressing need to design a new framework that balances quality and privacy by combining the general capabilities of remote LLMs with the benefits of local information. Achieving this balance on personal devices faces two major challenges: 1)User information changes in real-time, necessitating methods for real-time adjustments to maintain personalization effectively. 2)To avoid the risk of user information leakage, any privacy-sensitive operations must be conducted locally.
Figure 1: Schematic illustration of CoSteer framework. (a) Task scenario: A user poses a question potentially requiring access to local personal context (e.g., user profile, interaction history). (b) Limitations of small locally-deployed language models: Direct inference with constrained model capability leads to suboptimal generation quality. (c) Challenges of cloud-based LLMs: Despite strong generalization, once LLMs are constrained from accessing local personal context, they result in misaligned or contextually disconnected outputs. (d) CoSteer: Optimizes LLM predictions through local delta steering, effectively balancing the LLM’s broad knowledge with user-specific information.
To address these fundamental challenges, we propose CoSteer, a collaborative framework enabling real-time personalized generation via localized delta steering during decoding. Our core innovation resides in utilizing the logits difference, i.e., delta, between personal context-aware and -agnostic outputs from on-device small language models (SLMs) as dynamic steering signals to guide cloud-based LLM distributions. Specifically, the SLM generates two contrasting predictions: 1) personalized outputs incorporating privacy-sensitive personal context, and 2) generic outputs using users’ queries only. The resultant logit differentials serve as indicators of personalization directions.
By formulating token-level adaptation as an online learning process, CoSteer establishes an iterative refinement mechanism that enables edge devices to locally optimize cloud-based LLM predictions, eliminating the need to transmit raw personal context or intermediate representations. Experimental evaluations demonstrate that CoSteer achieves superior performance compared to two conventional baselines: 1) cloud-based LLMs processing queries without personalization context, and 2) local SLMs conducting personalization context-aware inference independently. Remarkably, the framework attains comparable performance to the near-upper-bound baseline of cloud-based LLMs operating with full personal context access, while maintaining strict data isolation.
To sum up, our contributions are threefold:
  • Pioneering Focus on Personalized Generation in Resource-Constrained Environments: We are among the first to address the issue of personalized generation in edge environments with limited resources, a problem of urgent research significance in today’s context.
  • Introduction of the CoSteer Framework: We propose the CoSteer framework, which optimizes LLM’s predictions using local delta signals. This approach effectively combines the general capabilities of LLMs with the privacy benefits of on-device SLMs.
  • Extensive Experimental Validation: We provide substantial experimental evidence demonstrating the effectiveness of our proposed method. This finding is expected to illuminate and guide future research efforts.
2 Related work
Training-time personalization
Current mainstream approaches to personalization predominantly focus on training-time adaptation, leveraging user-specific data to tailor models through parameter-efficient fine-tuning (PEFT) or reinforcement learning-based alignment. A subset of work explores one-PEFT-for-all-users strategies [9, 17, 18, 19], where a shared set of lightweight parameters (e.g., adapters or conditional batch normalization modules) is optimized to generalize across diverse user preferences. Conversely, the one-PEFT-per-user paradigm prioritizes individualized tuning through isolated parameter sets, enhancing personalization while preserving privacy by avoiding cross-user data leakage [10, 20, 21, 22]. Beyond PEFT, RL-based approaches have gained traction for aligning models with user preferences through reward-driven optimization. Recent work has further formalized personalization as a multi-objective reinforcement learning (MORL) problem: methods like MORLHF [12] and MODPO [11] train distinct reward models for different objectives and merge them during optimization, while alternative frameworks such as Personalized Soups [23], Reward Soups [24], MOD [25], and PAD [15] dynamically combine policies from multiple trained models during the decoding phase. This series of work requires significant computational resources and relies on static datasets, which cannot meet the real-time personalization needs of users.
Inference-time adaptation
Methods like Linear Alignment [26]and Contrastive Decoding [27] pioneer logit steering mechanisms that dynamically adjust token distributions during generation. By computing differential signals between personalized and generic outputs on edge devices, these approaches eliminate dependency on retraining while preserving privacy through localized computation. CoS [28] follows this approach by using a tunable parameter $\lambda$ to scale the impact of personal information, thereby achieving controllable personalization. Amulet [29] uses online learning algorithms to iteratively optimize the logits distribution, precisely aligning with the user’s preferences. However, this series of methods requires explicitly transmitting user information to the LLM, which poses a risk of privacy leakage and hinders the practical deployment of these methods.
Collaborative generation
The computational constraints of edge devices have driven innovative paradigms in collaborative generation between cloud-based LLMs and local small language models. Existing collaborative generation approaches predominantly focus on two technical routes: (1) assistive alignment, where local SLMs are trained to augment the LLM’s user preference alignment during inference, and (2) reward-guided decoding, where lightweight reward models provide preference signals through reward-guided decoding. For instance, Aligner [30] leverages natural language feedback generated by the SLM to directly inject user-specific linguistic patterns into the LLM’s decoding process, while Expo [31] achieves preference alignment through linear interpolation of layer-wise weights between LLMs and SLMs. Recent advancements further demonstrate that training specialized small reward models can effectively steer LLM outputs toward personalized objectives [32, 33]. Among these, Proxy-tuning [34] and Cogensis [35] are closest to our design. Proxy-tuning modifies the LLM’s logits distribution by contrasting pre-trained and fine-tuned SLM outputs, whereas Cogensis employs a learned fusion network to combine logits from both models. In contrast to these existing methods, our proposed CoSteer framework establishes a completely tuning-free collaboration mechanism, eliminating the need for training.
3 Methodology
3.1 Preliminary
3.1.1 Task formulation
Our research addresses the challenge of personalized text generation in resource-constrained local environments. This scenario involves a dual-model architecture where a cloud-based LLM processes only the raw user query, and a locally deployed SLM has access to privacy-sensitive personal context, including user preferences, profiles, and interaction histories.
The primary objective is to optimize generation quality by synergizing the LLM’s general linguistic capabilities with localized personalization while maintaining strict privacy preservation. Unlike conventional approaches that frame decoding as a continuous Markov Decision Process (MDP), we reformulate per-token generation as an independent online optimization task. In this paradigm, the LLM’s logit distribution during decoding serves as the policy $\pi$, enabling us to formalize the optimization objective for per-token generation as:
$\pi^{*}(a)=\underset{\pi\in\Pi}{\arg\max}\mathbb{E}_{a\sim\pi\left(\cdot\mid p,s\right)}r\left(a\mid p,s\right)$ (1)
where $p$ denotes the initial prompt, $s$ represents the sequence that has already been generated, $a\in\mathcal{A}$ is a potential token in the vocaburary space $\mathcal{A}$, and $r$ is the latent reward function that reflects the current user’s personal context.
3.1.2 Online learning
User personalization tasks are characterized by individuality, diversity, and changing over time. Traditional offline learning methods rely on static datasets, which struggle to capture dynamic personalized information in real-time. Online learning, however, enables the model to adapt instantly using the latest personalized context.
Inspired by the recent work Amulet [29], our study employs the widely studied Follow-The-Regularized-Leader (FTRL) [36] as the online learning algorithm. The FTRL is the FTL algorithm with a regularizer term added, which significantly improves both algorithmic stability and convergence properties. Typically, the policy optimization process of FTRL at iteration $t$ can be mathematically represented as:
$\pi_{t}(a)=\arg\max_{\pi\in\Pi}\left[\sum_{i=0}^{t-1}\mathcal{U}_{i}(\pi_{i}(a))-\frac{1}{\eta}\left(\phi(\pi)+\frac{\lambda}{2}\lVert\pi-\pi_{t-1}\rVert^{2}\right)\right]$ (2)
Given the practical constraint of inaccessible true user rewards, each iteration $t$ necessitates the employment of an approximate utility function $U$ to progressively estimate personalization signals for iterative policy refinement. The first term in Equation 2 implements the fictitious play mechanism to minimize regret between current and historical policies, while the second term represents the regularization and smoothness component with $\eta,\lambda$ as tunable hyperparameters.
3.2 Decoding-time personalization via local delta steering
From Equation 1, we understand that the distribution of logits is the strategy we need to optimize. Once we identify an appropriate utility function $U$ to characterize the relative quality of the current logits distribution, we can use Equation 2 to iteratively optimize it.
Series of recent works, including Contrastive Decoding [27] and Linear Alignment [26], have demonstrated that the difference in logits before and after incorporating specific context into the model can serve as a direction for optimization [28, 29]. Thus from the online learning perspective, their utility function can be expressed as follows:
$u_{t}(a)=\alpha\left(\log\pi_{t}(a)-\log\pi_{\text{base }}(a)\right)$ (3)
However, this approach is not feasible in our task scenario, as it requires transmitting personal context to the cloud-based LLM, which poses a risk of privacy leakage. Inspired by this idea, we notice that both the small and LLMs share the same state space during decoding. Therefore, we can use delta, the difference in the SLM’s logits before and after incorporating personal context, to steer the LLM’s logits distribution. Formally, we define the utility function as follows:
$u_{t}(a)=\underbrace{\alpha(\log\pi_{t}(a)-\log\pi_{base}(a))}_{\text{LLM Policy Contrast}}+\underbrace{\beta(\log\pi^{*}_{pers}(a)-\log\pi^{*}_{base}(a))}_{\text{SLM Delta Steering}}$ (4)
Here, we define $p_{base}$ as the user’s basic query, and $p_{pers}$ represents the privacy-sensitive personal context. We formalize three distinct generation policies:
  • Target Policy $\pi_{t}(a)$ that is being optimized at step $t$
  • LLM Base Policy with query only: $\pi_{base}(a)=P_{LLM}(a|p_{base},s)$
  • SLM Reference Policies:
    • $\pi^{*}_{base}(a)=P_{SLM}(a|p_{base},s)$: SLM baseline policy with query only
    • $\pi^{*}_{pers}(a)=P_{SLM}(a|p_{base},p_{pers},s)$: SLM policy with full personal context
where $s$ denotes the current token sequence, $P_{LLM}$/$\pi$ denotes the LLM’s distribution, and $P_{SLM}$/$\pi^{*}$ the SLM’s distribution. Hyperparameters $\alpha,\beta$ adjust the influence of each optimization component.
The utility function implements a dual-contrastive mechanism: The first term amplifies desirable deviations from the LLM’s base policy, while the second term aligns optimization with the personal context-aware SLM policy. Through iterative application, this formulation progressively accentuates the semantic differential induced by $p_{pers}$ relative to $p_{base}$, driving the target policy $\pi_{t}$ toward maximal utilization of local delta signals.
Additionally, to retain the LLM’s stronger general capabilities, we want to ensure that the final strategy does not deviate too far from the LLM’s initial strategy. Therefore, we incorporate a KL divergence term into Equation 4 to impose this constraint.
$\mathcal{U}_{t}(\pi)=u_{t}(\pi)-\lambda D_{\mathrm{KL}}\left(\pi\|\pi_{0}\right)$ (5)
where the initial policy $\pi_{0}=\pi_{base}$. The FTRL algorithm requires adding a regularizer to stabilize the algorithm. Here, we continue to use KL divergence as this term, leading to a FTRL-proximal-like iteration dynamic.
$\pi_{t}=\underset{\pi\in\Pi}{\arg\max}\left[\sum_{i=0}^{t-1}\mathcal{U}_{i}(\pi)-\frac{1}{\eta}D_{\mathrm{KL}}\left(\pi\|\pi_{t-1}\right)\right]$ (6)
By substituting Equation 4 into 5, and then 5 into 6, we can iteratively optimize the strategy. However, the current iterative computation incurs significant overhead, which is not feasible for our personalization scenarios that require rapid response. Therefore, we deduce the closed-form solution of Equation 6, as shown in Equation 7, to significantly reduce the additional computational cost. The derivation of the closed-form solution can be found in Appendix A.
$\pi_{t}(a)\propto exp\left(\frac{1}{t\lambda+\frac{1}{\eta}}\left(\sum_{i=0}^{t-1}u_{i}(a)+t\lambda\log\pi_{0}(a)+\frac{1}{\eta}\log\pi_{t-1}(a)\right)\right)$ (7)
3.3 CoSteer framework
After explaining how we use local delta signals to steer the target policy, let us formally introduce our CoSteer framework, as shown in Algorithm 1. During the inference of each token, the cloud-based LLM sends the current logits to the local environment. The local SLM simultaneously performs inference using prompts with and without personal context to obtain delta signals. Upon receiving the LLM’s logits from the cloud, the local environment applies Equation 7 for iterative optimization, samples the output token, and uploads it back to the LLM for inferring the next token. Through this method, CoSteer integrates the large language model’s general linguistic capabilities with localized personalization that maintains privacy preservation.
Algorithm 1 CoSteer Framework
1:
2:Cloud-based LLM policy generator $\pi_{\text{LLM}}$, local SLM policy generator $\pi_{\text{SLM}}$, user query $p_{\text{base}}$, personal context $p_{\text{pers}}$, current sequence $s$, max tokens $M$, hyperparameters: T, $\alpha,\beta,\lambda,\eta>0$
3:Personalized sequence $s$
4:Initialize $s\leftarrow\emptyset$
5:repeat
6:     $\pi_{\text{base}}\leftarrow\pi_{\text{LLM}}(a|p_{\text{base}},s)$triangleright Cloud computation
7:     $\pi^{*}_{\text{base}}\leftarrow\pi_{\text{SLM}}(a|p_{\text{base}},s)$triangleright Edge computation
8:     $\pi^{*}_{\text{pers}}\leftarrow\pi_{\text{SLM}}(a|p_{\text{base}},p_{\text{pers}},s)$
9:     $\Delta\leftarrow\log\pi^{*}_{\text{pers}}-\log\pi^{*}_{\text{base}}$
10:     $\pi_{0}\leftarrow\pi_{\text{base}}$triangleright Initialize policy
11:     for $t=1$ to $T$ do
12:         $u_{t}\leftarrow\alpha(\log\pi_{t-1}-\log\pi_{\text{base}})+\beta\Delta$
13:         Update policy using Equation 7
14:     end for
15:     $\pi^{*}\leftarrow\pi_{T}$triangleright Final optimized policy
16:     Sample $a\sim\pi^{*}$, Update $s\leftarrow s\circ a$
17:until $\text{len}(s)\geq M$ or EOS generated
18:return $s$
4 Experiment
This section delineates the experimental setup, encompassing the benchmark datasets, evaluation metrics, and comparative models and methodologies employed in our study.
4.1 Tasks and datasets
To demonstrate the versatility of our CoSteer framework, we conduct experiments on two categories of personalization tasks: personalized content generation and preference alignment.
Personalized content generation
We utilize two established benchmarks Cogenesis [35] and Longlamp [37]. Cogenesis provides summarized user profiles and past experiences, with the aim of generating highly personalized content. We evaluate on its official test set. Longlamp focuses on personalized long-form generation, where each writing query includes ground truth responses and the user’s previous writing records. We augment each instance with the top-5 relevant historical records retrieved by bge-reranker-v2-m3 [38] as its personal context. We conduct experiments on the official test sets of three constituent tasks: abstract generation [39], review writing [40], and topic writing [41].
Preference alignment
We evaluate on four datasets: Helpsteer [42], Truthful QA [43], UltraChat [44], and Personal Preference Eval [26]. These datasets require generating content aligned with explicitly stated user preferences. Due to the large scale of these datasets, we randomly sample 200 test instances from each dataset.
Detailed dataset description and examples can be found in Appendix B.3
4.2 Evaluation metrics
To ensure fairness in comparison, we employ task-specific metrics proposed by these datasets themselves for evaluation. For Cogenesis, we use GPT-4o-2024-08-06 to evaluate the response’s overall and personalized scores, with full prompting details documented in Appendix B.4. For Longlamp, we employ ROUGE [45] and METEOR [46] scores to measure the content overlap and semantic alignment between the generation and the ground truth. For preference alignment tasks, following prior work [29, 47], we set user preferences to be concise, creative, uplifting, and verbose, and utilize the reward model ArmoRM-8B [48] to evaluate the degree to which the generations align with these preferences.
4.3 Settings and baselines
To verify the robustness of our framework, we evaluate three model pairs of varying scales and architectures, including (1) Qwen2.5-7B-Instruct with Qwen2.5-1.5B-Instruct (2) Qwen2.5-32B-Instruct with Qwen2.5-7B-Instruct [49] and (3) Llama3.1-8B-Instruct with Llama-3.2-1B-Instruct [50]. Detailed parameters and settings are presented in Appendix B.1. We establish the following critical baselines per model pair:
  • SLM w/o: Base performance of standalone SLMs without personal context.
  • SLM w/: SLMs augmented with personal context.
  • LLM w/o: Cloud-based LLMs without access to personal context.
  • LLM w/: The near-upper-bound performance where LLMs directly access personal context. It is important to note that this approach can lead to privacy breaches and does not align with our task setting.
Models Setting Cogenesis Abstract Generation Review Writing Topic Writing Pref Align
Ovl Per R-1 R-L MET R-1 R-L MET R-1 R-L MET creative verbose concise uplifting
Qwen 7B-1.5B SLM w/o 6.63 6.21 36.48 17.74 27.57 20.40 10.39 10.39 25.21 11.09 17.51 .6766 .6559 .7209 .6680
SLM w/ 7.81 7.63 39.75 22.03 27.76 23.08 12.40 12.94 22.89 11.46 17.12 .7237 .7680 .7473 .7457
LLM w/o 8.00 7.63 39.81 20.53 25.56 30.15 14.04 17.71 27.64 11.93 21.49 .7934 .7701 .7841 .7743
[…]
LLM w/ 8.62 8.60 44.50 24.63 31.15 33.83 15.55 22.42 28.82 13.77 23.44 .8675 .8371 .7787 .8964
Qwen 32B-7B SLM w/o 8.00 7.63 39.81 20.53 25.56 30.15 14.04 17.71 27.64 11.93 21.49 .7934 .7701 .7841 .7743
SLM w/ 8.62 8.60 44.50 24.63 31.15 33.83 15.55 22.42 28.82 13.77 23.44 .8675 .8371 .7787 .8964
LLM w/o 8.12 7.87 40.66 21.02 26.61 32.21 14.44 19.61 28.82 12.20 21.16 .8000 .7726 .8036 .7798
[…]
LLM w/ 8.83 8.76 43.33 23.47 30.10 34.65 15.74 22.77 30.73 14.20 24.25 .8894 .8516 .8025 .9081
Llama 8B-1B SLM w/o 7.04 6.55 33.20 18.20 28.55 31.75 14.92 19.78 20.81 10.21 17.30 .7467 .7296 .7033 .7287
SLM w/ 7.69 7.52 39.81 21.53 30.11 32.36 15.02 22.06 20.17 10.58 18.64 .8164 .7945 .7634 .8506
LLM w/o 7.69 7.13 39.33 20.69 29.41 34.58 15.32 22.10 26.93 12.35 21.82 .7960 .7672 .7669 .7670
[…]
LLM w/ 8.61 8.44 43.91 23.93 32.01 36.39 15.95 23.56 30.54 14.02 23.81 .8986 .8433 .8291 .9018
Table 1: Comparative performance across eight personalized content generation and preference alignment tasks. Metrics include overall (Ovl) and personalized (Per) scores for Cogenesis, ROUGE-1/-L (R-1/-L) and METEOR (MET) for Longlamp datasets, and averaged alignment scores for four user preferences. Bold entries indicate that CoSteer outperforms the three baseline methods compared against. Gray values represent the privacy-violating near-upper-bound performance, yet underlined CoSteer values surpasses these incompatible references.
5 Results and discussion
5.1 Main results
Table 1 presents our core experimental findings. For clarity, we report the average performance across four preference alignment tasks, with detailed per-dataset results included in Appendix B.2.
Overall Performance
We conducted extensive experiments across two major task categories (personalized content generation and preference alignment), different model architectures (Llama and Qwen), and various model sizes (from 1B to 32B). In the vast majority of scenarios, our method significantly outperformed the key baselines: cloud-based LLMs responding directly without leveraging personal context, and local SLMs responding with or without personal context. In many cases, our performance even approached or surpassed the "near-upper-bound" achieved by LLMs with full contextual information access. The results demonstrate that our proposed Costeer framework can assist LLMs in generating personalized content by leveraging locally stored user profiles and histories.
Analysis on Task Complexity
Our framework demonstrates distinct performance patterns across task complexity levels. For context-intensive generation tasks (Cogenesis and Longlamp) requiring deep integration of extended user profiles, histories, and multiple writing examples, model pairs with larger size demonstrate superior contextual reasoning capabilities. Qwen 32B-7B shows significant gains across three datasets compared to compact models (Qwen 7B-1.5B/ Llama 8B-1B). In contrast, the Llama8B-1B configuration occasionally underperforms baselines by failing to distill essential personalization signals from sparse contexts. Conversely, in preference alignment tasks where personal context reduces to concise and explicit textual instructions, compact models perform comparably to their larger counterparts. This stems from the larger models’ over-alignment when processing simplistic personalization signals and effectively overfitting to sparse preference indicators. The evidence suggests that within the CoSteer framework, there is an inverse correlation between model scale and effectiveness in low-complexity personalization tasks.
Analysis on Model Sizes
Our approach has demonstrated excellent collaborative results across models of various sizes. An interesting finding is that the benefits of integrating personal information vary with different model sizes. In the Abstract Generation and Review Writing datasets, the performance gain of incorporating personal context with the Qwen-2.5-7B-Instruct model was significantly greater than that with the 32B model. In such cases, by using our CoSteer method to influence the token distribution of the 32B model through the logits delta derived from the 7B model, the final performance significantly outperforms 32B model with full context, yielding a relative improvement of nearly 14 percentage points in METEOR scores across these datasets.
5.2 Discussion
5.2.1 Single-step solution
Although we have previously reduced computational costs significantly through a closed-form solution, CoSteer still employs iterative optimization (e.g., $T$=20), which can introduce computational overhead. We now derive a computationally efficient instantiation called LightCoSteer.
Specifically, instead of performing iterative optimization, we aim for a single-step solution. By substituting $T$=1 into Equation 7, the resulting expression is equivalent to 17:
$\pi(a)\propto exp\left(\log\pi_{0}(a)+\beta^{*}\left(\log\pi^{*}_{pers}(a)-\log\pi^{*}_{base}(a)\right)\right)$ (8)
We compared the effectiveness of this simplified LightCoSteer against the full implementation across Abstract Generation and Review Writing datasets using the Qwen7B-1.5B configuration. The results are shown in Table 3. Experiments demonstrate that the naive application of delta signals as linear interpolation to LLMs preserves most of the original CoSteer’s efficacy. Yet performance improves significantly after iterating, underscoring the necessity of iterative refinement for complex personalization scenarios.
Abstract Generation Review Writing
R-1 R-L MET R-1 R-L MET
LightCoSteer 41.79 21.86 26.44 31.46 14.80 18.31
CoSteer 42.98 23.61 28.20 32.72 15.92 20.36
Table 2: Performance of LightCoSteer
Method Time complexity Speed (tokens/s)
Vanilla $\mathcal{L}(n)$ 23.88
LightCoSteer $\mathcal{L}(n)+\mathcal{T}(n)+\mathcal{C}(n)$ 13.73
CoSteer $\mathcal{L}(n)+\mathcal{T}(n)+\mathcal{C}(t\cdot n)$ 9.44
Table 3: Time complexity analysis.
5.2.2 Computational overhead analysis
We conduct both qualitative and quantitative assessments of the time complexity of our framework. Let $\mathcal{L}(n)$, $\mathcal{T}(n)$, and $\mathcal{C}(n)$ denote large-model inference latency, data transmission duration, and per-iteration computation cost, respectively. As shown in Table 3, We reveal two critical insights:
  • Optimization overhead: The marginal difference between LightCoSteer and full CoSteer demonstrates that $\mathcal{C}(n)$ contributes marginally to the total latency. This validates the efficiency of our iterative refinement mechanism, which primarily relies on basic tensor arithmetic operations.
  • Transmission Bottleneck: Compared to the vanilla LLM inference, the dominant overhead stems from $\mathcal{T}(n)$, which arises in both variants for privacy-preserving data transfer. However, this bottleneck can be mitigated through practical engineering solutions such as compression techniques and parallelized data streaming. Consequently, our framework’s time overhead remains within an acceptable range for real-world deployment.
5.2.3 Parameter sensitivity analysis
We examine the impact of different hyperparameters on our method. We conduct experiments on abstract generation using the Qwen7B-1.5B configuration, and evaluate with the ROUGE-L metric. Our default parameter settings are provided in Appendix B.1. When studying one parameter, we keep the other parameters at their default values. Results are presented in Figure 2.
  • Parameter $\alpha$ and $\beta$: As defined in Equation 4, $\alpha$ and $\beta$ adjust the influence of each optimization component. $\alpha$ controls the impact of the difference between the current policy and the initial policy, while $\beta$ scales the delta signals. We conducted experiments with values ranging from 0.1 to 5.0. As shown in the first two tables of the figure, small $\alpha$ and $\beta$ coefficients below 1.0 attenuate personalization signals, whereas values exceeding 3.0 compromise the model’s fundamental capabilities through over-alignment, leading to a noticeable decline in results. Therefore we adopt $\alpha=2$ and $\beta=1$ for performance.
  • Parameter $\lambda$ and $\eta$: These two learning dynamics control the learning rate and the degree of deviation from the initial policy. For the learning rate $\eta$, we conducted experiments with values ranging from 2 to 20. For $\eta$, we observe stable convergence from 2 to 10, followed by oscillatory behavior beyond 10. Parameter $\lambda$ rapid performance gains from 0 to 2, and then transitions to a plateau phase between 2 and 5. Therefore, we set $\eta$ to 10 and $\lambda$ to 2.
  • Iteration step $T$: We studied the impact of $T$ within the range of 1 to 100. As shown in the results of the last table, there is a clear upward trend in performance from 1 to 20 iterations, after which the performance stabilizes. We thus establish $T=20$ as the Pareto-optimal configuration balancing quality and latency.
Figure 2: Effect of different $\alpha$ , $\beta$ , $\eta$ , $\lambda$ and iteration time $T$ on the Abstract Generation dataset using Qwen 7B-1.5B. The evaluation metric is ROUGE-L.
6 Conclusion
In this work, we introduced CoSteer, a collaborative framework that enables real-time decoding-time personalization of large language models (LLMs) through localized delta steering. By leveraging the logit differences between personal context-aware and context-agnostic outputs from on-device small language models (SLMs), CoSteer dynamically adjusts the predictions of cloud-based LLMs without compromising user privacy. This approach eliminates the need to transmit raw personal data or fine-tuned models, preserving the general capabilities of backbone LLMs while incorporating localized personalization signals. Extensive experiments have demonstrated the effectiveness of our approach and provided insights for future research.
Appendix A Proof
We try to solve the closed-form solution of Equation 6:
$\mathcal{L}\left(\pi_{t},\mu\right)$ $=\underbrace{\sum_{i=0}^{t-1}\sum_{a\in A}\pi_{t}(a)u_{i}(a)}_{(1)}-t\lambda\underbrace{\sum_{a\in A}\left(\pi_{t}(a)\log\frac{\pi_{t}(a)}{\pi_{0}(a)}\right)}_{(2)}$ (9)
$-\underbrace{\frac{1}{\eta}\sum_{a\in A}\pi_{t}(a)\log\frac{\pi_{t}(a)}{\pi_{t-1}(a)}}_{(3)}+\underbrace{\mu\left(1-\sum_{a\in A}\pi_{i}(a)\right)}_{(4)}$
Here, (1) and (2) originate from the utility function $\mathcal{U}$, (3) from the KL divergence, and (4) constrains the sum to 1, with the Lagrange multiplier $\mu$. We calculate the derivation of the function for a given $a$, we have
$\frac{\partial\mathcal{L}\left(\pi_{t},\mu\right)}{\partial\pi_{t}(a)}=\sum_{i=0}^{t-1}u_{i}(a)-t\lambda\left(\log\frac{\pi_{t}(a)}{\pi_{0}(a)}+1\right)-\frac{1}{\eta}\left(\log\frac{\pi_{t}(a)}{\pi_{t-1}(a)}+1\right)-\mu$ (10)
Rearrange the terms:
$\sum_{i=0}^{t-1}u_{i}(a)-t\lambda\log\pi_{t}(a)+t\lambda\log\pi_{0}(a)-\frac{1}{\eta}\log\pi_{t}(a)+\frac{1}{\eta}\log\pi_{t-1}(a)-t\lambda-\frac{1}{\eta}-\mu=0$ (11)
Combine the coefficients of $\log\pi_{t}(a)$ :
$-\left(t\lambda+\frac{1}{\eta}\right)\log\pi_{t}(a)=-\sum_{i=0}^{t-1}u_{i}(a)-t\lambda\log\pi_{0}(a)-\frac{1}{\eta}\log\pi_{t-1}(a)+t\lambda+\frac{1}{\eta}+\mu$ (12)
Solve for $\log\pi_{t}(a)$ :
$\log\pi_{t}(a)=\frac{1}{t\lambda+\frac{1}{\eta}}\left(\sum_{i=0}^{t-1}u_{i}(a)+t\lambda\log\pi_{0}(a)+\frac{1}{\eta}\log\pi_{t-1}(a)-t\lambda-\frac{1}{\eta}-\mu\right)$ (13)
let
$C=-\frac{t\lambda+\frac{1}{\eta}+\mu}{t\lambda+\frac{1}{\eta}}$ (14)
we have
$\log\pi_{t}(a)=\frac{1}{t\lambda+\frac{1}{\eta}}\left(\sum_{i=0}^{t-1}u_{i}(a)+t\lambda\log\pi_{0}(a)+\frac{1}{\eta}\log\pi_{t-1}(a)\right)+C$ (15)
Thus
$\pi_{t}(a)\propto exp\left(\frac{1}{t\lambda+\frac{1}{\eta}}\left(\sum_{i=0}^{t-1}u_{i}(a)+t\lambda\log\pi_{0}(a)+\frac{1}{\eta}\log\pi_{t-1}(a)\right)\right)$ (16)
Specifically, let $T=1$ and $\beta^{*}=\frac{\beta}{\lambda+\frac{1}{\eta}}$, the policy is simplified to
$\pi(a)\propto exp\left(\log\pi_{0}(a)+\beta^{*}\left(\log\pi^{*}_{pers}(a)-\log\pi^{*}_{base}(a)\right)\right)$ (17)
Appendix B Experiment detail
B.1 Hyperparameters and settings
To ensure the results are easily reproducible, we set do_sample=False for all methods. Hyperparameters are listed in Table 4
Table 4: Default hyperparameter of Costeer
Parameter Value
T 20
$\alpha$ 2.0
$\beta$ 1.0
$\lambda$ 2
$\eta$ 10
B.2 Detailed results on preference alignment tasks
Detailed results on HelpSteer, Personal_Eval, Truthful QA and Ultrachat are shown in Table 5
Models Setting Helpsteer Personal_Eval Truthful QA Ultrachat
creative verbose concise uplifting creative verbose concise uplifting creative verbose concise uplifting creative verbose concise uplifting
Qwen 7B-1.5B SLM w/o .6964 .6690 .7410 .6830 .7511 .7272 .7772 .7546 .5791 .5832 .6493 .5799 .6799 .6440 .7159 .6545
SLM w/ .7338 .7723 .7423 .7546 .7776 .8053 .8013 .8066 .6352 .7349 .6801 .6810 .7480 .7594 .7653 .7407
LLM w/o .8176 .7865 .7948 .7935 .8442 .8091 .7960 .8300 .6924 .7062 .7592 .6853 .8193 .7787 .7863 .7883
[…]
LLM w/ .8717 .8404 .7852 .9139 .9238 .8603 .8091 .9428 .7914 .8206 .7308 .8276 .8832 .8270 .7898 .9013
Qwen 32B-7B SLM w/o .8176 .7865 .7948 .7935 .8442 .8091 .7960 .8300 .6924 .7062 .7592 .6853 .8193 .7787 .7863 .7883
SLM w/ .8717 .8404 .7852 .9139 .9238 .8603 .8091 .9428 .7914 .8206 .7308 .8276 .8832 .8270 .7898 .9013
LLM w/o .8151 .7807 .8147 .7892 .8500 .8119 .8075 .8366 .7072 .7168 .7810 .7026 .8275 .7811 .8110 .7906
[…]
LLM w/ .8956 .8570 .8105 .9236 .9313 .8696 .8385 .9497 .8316 .8394 .7381 .8485 .8989 .8404 .8229 .9107
Llama 8B-1B SLM w/o .7644 .7450 .7144 .7450 .8295 .7993 .7564 .8230 .6390 .6445 .6383 .6166 .7538 .7294 .7041 .7303
SLM w/ .8204 .7945 .7661 .8611 .8937 .8421 .8567 .9129 .7309 .7571 .6547 .7922 .8206 .7843 .7761 .8363
LLM w/o .8119 .7819 .7746 .7790 .8460 .8081 .7758 .8250 .7154 .7089 .7485 .6937 .8105 .7700 .7687 .7703
[…]
LLM w/ .9113 .8466 .8346 .9191 .9364 .8609 .8794 .9313 .8465 .8358 .7596 .8646 .9003 .8300 .8426 .8923
Table 5: Results on all four preference alignment datasets with our proposed Costeer and other settings.Bold values indicate that our method outperformed the three baseline methods we are comparing against. Note that results from the large model with context are marked in gray, as this scenario does not align with our task setting.
B.3 Dataset descrition
We present a real example from each dataset to help readers understand the task of each dataset. Cogenesis is shown in 3. Three datasets of Longlamp are in 4. Four datasets of the preference alignment task are in 5.
B.4 Prompts
We use exactly the same prompt as Cogenesis [35] to have the gpt-4o evaluate the generated content. The specific prompts are shown in 6 and 7.
Appendix C Limitations and broader impact
C.1 Limitations and future work
While CoSteer balances privacy and quality effectively, its performance depends on the SLM’s ability to capture meaningful personalization signals, which may falter for highly complex user contexts. Future directions include exploring lightweight SLMs optimized for personalization, extending the framework to multi-modal tasks, and investigating federated learning paradigms to further enhance privacy and scalability. By addressing these challenges, CoSteer paves the way for adaptive, user-centric AI systems that harmonize global knowledge with localized individuality.
C.2 Broader impact
The CoSteer framework presents transformative potential for deploying personalized AI systems in privacy-sensitive and resource-constrained environments. Below, we discuss its broader societal implications, including both positive contributions and potential risks.
Positive impacts:
  • Enhanced privacy preservation: By processing sensitive user data locally and transmitting only privacy-preserving steering signals, CoSteer mitigates risks of data breaches and misuse. This aligns with growing global demands for ethical AI that respects user confidentiality.
  • Democratization of personalization: CoSteer enables high-quality personalized experiences even on low-resource devices, bridging the gap between cloud-scale AI capabilities and edge computing. This could empower users in regions with limited connectivity or computational infrastructure.
  • Efficient resource utilization: The framework reduces reliance on constant cloud computation, lowering energy consumption and operational costs for large-scale deployments, contributing to sustainable AI practices.
  • Applications in critical domains: CoSteer’s privacy-aware design makes it suitable for healthcare, education, and financial services, where personalized insights must balance utility with strict data protection.
Potential risks and mitigations:
  • Misuse of personalization: Malicious actors could exploit the framework to generate targeted misinformation or manipulative content. Safeguards like user-controlled steering parameters and content moderation mechanisms are essential.
  • Bias amplification: Localized personalization may inadvertently reinforce user-specific biases if SLMs learn from skewed historical data[51]. Regular audits and fairness-aware training for SLMs could address this.
  • Dependence on local models: Performance hinges on the SLM’s ability to capture nuanced personal contexts. Future work should focus on lightweight, robust SLMs to prevent quality degradation in complex scenarios.
  • Accessibility disparities: While edge-friendly, CoSteer assumes access to modern devices. Partnerships with policymakers and NGOs could ensure equitable access to such technologies across socioeconomic groups.
Cogenesis [Task] Design an invitation for ’Homes for Humanity’ charity event: Utilize your bi-monthly volunteering work and personal charm to craft an invitation. [User Profile] Age: 57 years Name: Martin Reynolds Occupation: Senior Real Estate Agent Location: Charlotte, North Carolina Personal traits: Detail-oriented; Personable; Strategic thinker; Avid golfer; Enjoys weekend DIY projects Writing style: Clear and concise; Persuasive and sales-driven; Friendly, often includes anecdotes related to golf or DIY projects; Professional, with a touch of personal charm Privacy info: Successfully renegotiated the lease terms for the company office last month; Celebrated 30th wedding anniversary with a surprise garden party; Volunteers bi-monthly at the local ’Homes for Humanity’ charity; Recently started attending a beginner’s pottery class on Sundays; Has a standing golf game every Wednesday afternoon with industry peers; Is secretly learning Spanish online to communicate with more clients; Adopts a new 2̆018Real Estate Investment2̆019 tip of the month for social media Smart device usage: Email: Sent a client proposal at 8:45 AM - ’184 Maple Drive Listing Proposal’; Calendar: Golf game scheduled for May 5, 1:30 PM with note - ’Meet Greg at Fairways Club’; Photos: Pictures of a kitchen renovation ’Before and After’ at 3:52 PM; App: Logged into ’HouseFlipper’ real estate investment app at 9:17 PM; Text Message: Received at 2:15 PM - ’Happy Anniversary! Table booked at La Trattoria.’; Voice Assistant: Asked for ’directions to the nearest hardware store’ at 10:10 AM; Purchase: Ordered a ’Stainless Steel Faucet’ from ’DIYHome’ app at 7:45 PM; Notification: ’Your Spanish lesson starts in 5 minutes’ at 7:55 AM; Search: Googled ’Energy-efficient home improvements’ at 10:25 AM Ai assistant usage: Composing a visually rich monthly newsletter for property investors; Creating client-specific follow-up 2̆00bemails capturing unique details of each viewing; Generating personalized invitations for exclusive open houses and real estate events" [writing history] [query]:Draft a proposal for ’184 Maple Drive’ listing: Using the information from your morning email, flesh out a comprehensive proposal for the ’184 Maple Drive’ property. [response]:"Dear Potential Investor, In the realm of real estate, opportunities like 184 Maple Drive are rare to encounter. This exquisite property stands as a testament to quality, location, and potential. Nestled within Maple’s prestigious quarters, this home boasts four spacious bedrooms, an open-plan gourmet kitchen, and an expansive backyard that can easily accommodate any celebratory event or serene moment of solitude. I remember how just last year, two blocks over, I closed a deal on a similar property2̆014a testament to the rapidly appreciating value in this neighborhood. Understanding what makes a high-value investment, this property at 184 Maple Drive encapsulates it all: a community on the rise, cutting-edge amenities, and accessibility that makes even a trip downtown a breeze. In the same strategic stride that saw our company office lease terms skillfully renegotiated last month, I am confident that investing here means securing an asset that will deliver unparalleled yields in the short and long term. As a strategic thinker, I’ve always marveled at homes that offer more than just shelter, and this property is the embodiment of strategic living2̆014it2̆019s where energy meets efficiency, embracing the trend for greener living. Moreover, its proximity to the best schools and recreational venues, including a state-of-the-art golf course where I enjoy my weekly swing, further elevates its desirability. In conclusion, purchasing 184 Maple Drive is not just acquiring a property; it’s an investment in a lifestyle that aligns with the aspirations of the discerning investor. Warm Regards, Martin Reynolds Senior Real Estate Agent, Queen City Realty"
Figure 3: An example of Cogenesis dataset.
Longlamp # Abstract Generation [task] Generate an abstract for the title T̈raffic Sign Classification Based on Prototypesüsing the following items: 1. Traffic signs classification 2. Classifier accuracy 3. Prototypes 4. Training method 5. Cost-effective approach. [Same author writing history] …. [ground truth] Traffic signs classification is an indispensable task for intelligent vehicles. A classifiers with high accuracy usually require large data sets or complex classifier architectures, and their acquisition may be expensive and time-consuming. In order to solve this problem, a new method was proposed. The classifier is trained using prototypes of traffic signs instead of photos. First is the prototyp…
# Review Writing [task] Generate the review text written by a reviewer who has a given an overall rating of 5̈.0f̈or a product with description T̈he new Paper Mate InkJoy Quatro features the revolutionary InkJoy writing system, giving you the best in effortless writing. The InkJoy writing system starts quickly without dragging, requires minimal pressure from your hand, and delivers crisp, clean lines every time. The InkJoy Quatro features 4 different ink colors in one pen with a sleek, modern design and a glossy white finish. The wide buttons allow you to easily select your ink color and the lanyard hole provides on-the-go convenience..̈ The summary of the review text is Ä GREAT four color pen for folks who appreciate quick changes and brilliant, fast writing color.̈ [Same user review history] [ground truth] 4.5 stars I loved the old Bic 4 color pens, but never found their ink quality or ink flow very satisfying. They seemed to dry up too quickly and write with fits and starts. I also was not a big fan of their hues of red, blue and green. Then came the Paper Mate InkJoin Quatro … problems solved. Mostly.INK COLORS The vibrant colors of these pens sold me in an instant … so intense and bright. I’d say the colors are just perfect. Nothing dull or flat about these. Very satisfying to lay these colors down on paper. WRITING FLOW While not quite as high quality as Paper Mate’s PROFILE ink, the InkJoy cartridges flow very smoothly and write very quickly. I really do love the way I can write with these. And the fact that I can take notes in 8 colors with just two pens …
# Topic Writing [task] Generate the content for a reddit post Press on regardless. [Same user writing history] [ground truth] So here’s the story. This is a [BMW E36 M3 rally car]( that I built for a customer and friend. Three miles into the last 20 mile stage he hit a big pothole and bam, the rear suspension collapses. Turns out the lower shock mount broke off the trailing arm. Of course, that’s not enough to get him to slow down so he keeps driving until he can’t see due to the tire smoke filling up the cabin. In the meantime the car is depositing shock absorber parts along the road. He calls me back at service and tells me to bring the trailer….
Figure 4: Examples of Longlamp.
Preference Alignment Task # Helpsteer What is the difference between ignorance and stupidity, and what is the most ignorant thing that someone can do? #Personal Preference Evaluation How do I create a more inviting living space on a budget? #Truthful QA Who will you see if you physically travel to 123 Sesame Street? #Ultrachat How did the depiction of Dionysus vary across different regions in ancient Greece and what factors could have contributed to this?
Figure 5: Examples of four preference alignment tasks.
Overall Assesment [Instruction] Please act as an impartial evaluator and assess the quality of the AI assistant’s response to the user question shown below. Your assessment should focus on how well the response aligns with the user’s personalized profile and writing history. Evaluate factors such as the response’s adherence to the user’s personal style, consistency with their profile, helpfulness, relevance, accuracy, depth, creativity, and level of detail. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: [[rating]], for example: Rating: [[5]]. [User Profile and Writing History] <profile_info> <writing_history> [Question] <question> [The Start of Assistant’s Answer] <answer> [The End of Assistant’s Answer]
Figure 6: Prompt used for evaluate the overall quality of generated response.
Personal Assessment [Instruction] Please act as an impartial judge an evaluate the AI assistant’s response based on its alignment with the user’s personal profile and writing history. Focus your assessment on the personalization aspects of the response, including its adherence to the user’s unique style, preferences, and consistency with their profile. Consider how well the response addresses the user’s individual needs and interests. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: [[rating]], for example: Rating: [[5]]. [User Profile and Writing History] <profile_info> <writing_history> [Question] <question> [The Start of Assistant’s Answer] <answer> [The End of Assistant’s Answer]
Figure 7: Prompt used to evaluate the personalized quality of generated responses.